### v1

# Manuscript submitted to Aequitas 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2024, Santiago de Compostela, Spain

Abstract
Motivated by the recital (67) of the current corrigendum of the AI Act in the European Union, we propose and present measures and mitigation strategies for discrimination in tabular datasets. We specifically focus on datasets that contain multiple protected attributes, such as nationality, age, and sex. This makes measuring and mitigating bias more challenging, as many existing methods are designed for a single protected attribute. This paper comes with a twofold contribution: Firstly, new discrimination measures are introduced. These measures are categorized in our framework along with existing ones, guiding researchers and practitioners in choosing the right measure to assess the fairness of the underlying dataset. Secondly, a novel application of an existing bias mitigation method, FairDo, is presented. We show that this strategy can mitigate any type of discrimination, including intersectional discrimination, by transforming the dataset. By conducting experiments on real-world datasets (Adult, Bank, Compas), we demonstrate that de-biasing datasets with multiple protected attributes is achievable. Further, the transformed fair datasets do not compromise any of the tested machine learning models’ performances significantly when trained on these datasets compared to the original datasets. Discrimination was reduced by up to 83% in our experimentation. For most experiments, the disparity between protected groups was reduced by at least 7% and 27% on average. Generally, the findings show that the mitigation strategy used is effective, and this study contributes to the ongoing discussion on the implementation of the European Union’s AI Act.
keywords:
[…]
1 Introduction
Discrimination in artificial intelligence (AI) applications is a growing concern since the adoption of the AI Act by the European Parliament on March 13, 2024 [1]. It still remains a significant challenge across numerous domains [2, 3, 4, 5]. To prevent biased outcomes, pre-processing methods are often used to mitigate biases in datasets before training machine learning models [6, 7, 8, 9]. These methods are even mentioned in Recital (67) at the current corrigendum of the AI Act [1]:
“[…] The data sets should also have the appropriate statistical properties, including as regards the persons or groups of persons in relation to whom the high-risk AI system is intended to be used, with specific attention to the mitigation of possible biases in the data sets […]”
Since datasets often consist of multiple protected attributes, pre-processing methods should be able to handle these cases. However, only a few works have addressed it  [7, 10, 11, 12] and de-biasing datasets with multiple protected attributes is still an ongoing research topic.
Our paper mainly focuses on how to measure and mitigate discrimination in datasets where multiple protected attributes are present. In our first contribution, we provide a comprehensive categorization of discrimination measuring methods. Besides introducing new measures for some of these cases, we also categorize existing measures from the literature. Some of the listed measures specifically address intersectional discrimination and non-binary groups. The second contribution deals with bias mitigation. For this, we use an existing pre-processing framework, FairDo [9], that is fairness-agnostic. The fairness-agnostic property makes it possible to define any discrimination measure that should be minimized. By implementing the introduced measures, we can therefore mitigate biases for multiple protected attributes. Another advantage of FairDo is that it preserves data integrity and does not modify the features of individuals during the optimization process, unlike other methods [13, 3, 7].
We evaluated our methodology on popular tabular datasets with fairness concerns, such as Adult [14], Bank [15], and Compas [16]. These datasets contain multiple protected attributes and we mitigated the inherent biases. We used different discrimination measures to evaluate the effectiveness of the bias mitigation process. Because a successful mitigation process does not guarantee that the outcomes of machine learning models are fair, we trained machine learning models on the transformed datasets and evaluated their predictions regarding fairness and performance.
The results of the bias mitigation process as well as the performance of the machine learning models are promising. They indicate that achieving fairness in datasets with multiple protected attributes is perhaps possible, and FairDo is a proper framework for this task. Overall, our work contributes technical solutions for stakeholders to enhance the fairness of datasets and machine learning models, aiming for compliance with the AI Act [1].
Figure 1: Stick figures can be differentiated by their color and shape. In intersectional discrimination, attributes are intersected, which leads to new subgroups. In non-intersectional, each attribute is treated separately, i.e., colors and shapes are not intersecting in this case.
2 Preliminaries
To handle multiple protected attributes, we define $Z=\{Z_{1},\ldots,Z_{p}\}$ as a set of protected attributes. It can represent the set of sociodemographic features such as age, gender, and ethnicity. These factors may make individuals vulnerable to discrimination. Each protected attribute $Z_{k}\in Z$ is formally a discrete random variable that can take on values from a set $g_{k}$. In this context, we refer $g_{k}$ to groups that describe distinct social categories of a protected attribute. For example, let $Z_{k}$ represent gender; then $g_{k}$ is a set containing the genders male, female, and non-binary. To avoid limitations to a particular group fairness notion, we introduce a generalized notation based on the works of Žliobaitė [2], Duong and Conrad [9] in the following.
Definition 2.1 (Treatment).
Let $E_{1},E_{2}$ be events and $Z_{k}$ be a random variable that can take on values from $g_{k}$, then we call the conditional probability
$P(E_{1}\mid E_{2},Z_{k}=i)$
treatment, where $i\in g_{k}$. $E_{1}$ describes some favorable outcome, such as getting accepted for a job, while $E_{2}$ often represents some additional information about the individual, such as their qualifications.
Definition 2.2 (Fairness Criteria).
With the definition of treatment, we can define fairness criteria that demand equal treatments for different groups. Let $P(E_{1}\mid E_{2},Z_{k}=i)$ and $P(E_{1}\mid E_{2},Z_{k}=j)$ be two treatments, then we call the following equation:
$P(E_{1}\mid E_{2},Z_{k}=i)=P(E_{1}\mid E_{2},Z_{k}=j)$
a fairness criterion, for all $i,j\in g_{k}$.
Definition 2.2 allows us to define various group fairness criteria, including statistical parity [17], predictive parity [3], equality of opportunity [18], etc. They all demand some sort of equal outcome for different groups and can be defined by configuring the events $E_{1},E_{2}$. For instance, statistical parity [17] requires that two different groups have an equal probability of receiving a favorable outcome $(Y=1)$.
Example 2.1 (Statistical Parity [17]).
To define statistical parity for the attribute $Z_{k}$ using our notation, we set $E_{1}\coloneqq(Y=1)$ and $E_{2}\coloneqq\text{True}$. By setting $E_{2}$ to a tautology, we compare the probabilities of the event $Y=1$ across different groups without conditioning on any additional event:
$P(Y=1\mid\text{True},Z_{k}=i)$ $=P(Y=1\mid\text{True},Z_{k}=j)$
$\iff P(Y=1\mid Z_{k}=i)$ $=P(Y=1\mid Z_{k}=j),$
where $i,j\in g$ represent different groups.
In real-world applications, achieving equal probabilities for certain outcomes is not always possible. Due to variations in sample sizes in the groups, it is common to yield unequal treatments, even when they are similar. Thus, existing literature [2] uses the absolute difference to quantify the strength of discrimination.
Definition 2.3 (Disparity).
Let $P(E_{1}\mid E_{2},Z_{k}=i)$ and $P(E_{1}\mid E_{2},Z_{k}=j)$ be two treatments, then we refer to
$\delta_{Z_{k}}(i,j,E_{1},E_{2})=|P(E_{1}\mid E_{2},Z_{k}=i)-P(E_{1}\mid E_{2},Z_{k}=j)|$
as the disparity, for all $i,j\in g_{k}$. Trivially, $\delta_{Z_{k}}$ is commutative regarding $i,j$. In practice, it prevents reverse discrimination due to the absolute value.
Definition 2.4 (Discrimination).
We use $\psi\colon\mathbb{D}\rightarrow\mathbb{R}$ to denote some discrimination measure that quantifies the discrimination inherent in any dataset $\mathcal{D}\in\mathbb{D}$. A dataset $\mathcal{D}$ consists of features, protected attributes, and labels for each individual. The explicit form of $\psi$ depends on cases introduced in Section 3.
3 Measuring Discrimination for Multiple Attributes
We found that numerous scenarios arise when dealing with multiple protected attributes. We categorize these scenarios based on the number of groups, denoted as $|g|$, and number of protected attributes, denoted as $|Z|$. By going through all cases, we present possible approaches from the literature as well as our own suggestions to measure discrimination.
3.1 Single Protected Attribute $(|Z|=1)$
In case of having only one protected attribute, i.e., $|Z|=|\{Z_{1}\}|=1$, we distinguish between cases by the number of available groups $|g|$ in the dataset. We categorize the cases by $|g|=0,1,2$, and $|g|>2$.
3.1.1 No Groups ($|g|=0$)
When having no groups, the measurement of discrimination is impossible if no assumptions are being made. Discrimination can be assessed through proxy variables [19]; however, this approach can be imprecise and may introduce new biases. This case is equivalent to having no protected attribute, i.e., $|Z|=0$.
3.1.2 Single Group ($|g|=1$)
Similarly to the case of having no groups, discrimination cannot be measured when having only one group. For this, we propose practices where prior information can be incorporated:
  1. 1.
    No discrimination: As no difference towards any other group can be measured, returning a discrimination score of 0 is one viable option.
    $\psi(\mathcal{D})=0.$ (1)
  2. 2.
    Difference to optimal treatment: Another way is to return the absolute difference of the group’s outcome to the optimal treatment. For example, group $i$ has an 80% chance of getting the favorable treatment. Ideally, having a 100% chance of receiving the favorable treatment would represent the optimal scenario. Therefore, the discrimination score is 20% in this case. It is given by:
    $\psi(\mathcal{D})=|P(E_{1}\mid E_{2},Z_{1}=i)-1|.$ (2)
  3. 3.
    Difference to expected treatment: We can use the expected treatment as a reference point. For example, we know that a company has a 50% acceptance rate for job applications. Now a machine learning classifier is trained to predict whether an applicant will be accepted and the model’s predictions result in a 60% acceptance rate for group $i$. Hence, the model is positively biased towards group $i$ by 10%. This can be formulated as:
    $\psi(\mathcal{D})=|P(E_{1}\mid E_{2},Z_{1}=i)-p_{\text{expect.}}|,$ (3)
    where $p_{\text{expect.}}$ is the expected treatment. It can describe the average treatment across all groups [20] or some other prior information that is not included in the dataset.
3.1.3 Binary Groups ($|g|=2$)
Without using any prior information, we can calculate the discrimination score by taking the absolute difference between the treatments of the two groups, as advised by Žliobaitė [2]. The discrimination measure $\psi$ is then simply given by the disparity as mentioned in Definition (2.3).
3.1.4 Non-binary Groups ($|g|>2$)
While the case for binary attributes is straightforward, it becomes non-trivial for non-binary attributes that arise naturally in real-world data. We can fall back to $|g|=2$ by calculating the absolute difference between every distinct group $i,j\in g$. Because the discrimination between $i$ and $j$ is the same as between $j$ and $i$, only $\binom{|g|}{2}$ pairs need to be compared and we use an aggregation function $\operatorname{agg}^{(1)}$ to report the differences [2]. Lum et al. [21] refer to measures that aggregate or summarize discrimination scores as meta-metrics. The aggregate can be the sum or maximum function, depending on the use case. The result for a single protected attribute $Z_{k}$ with two or more groups can be computed as follows:
$\psi(\mathcal{D})=\underset{i,j\in g_{k},i<j}{\operatorname{agg}^{(1)}}\delta_{Z_{k}}(i,j,E_{1},E_{2}),$ (4)
where $\delta_{Z_{k}}$ is the disparity as defined in Definition (2.3) and $i<j$ ensures that each pair is considered only once (assuming label encoded groups). According to Žliobaitė [2] and her personal discussions with legal experts, she advocates using the maximum function, i.e.,
$\psi(\mathcal{D})$ $=\underset{i,j\in g_{k},i<j}{\max}\delta_{Z_{k}}(i,j,E_{1},E_{2})$ (5)
$=\underset{i\in g_{k}}{\max}\,P(E_{1}\mid E_{2},Z_{k}=i)-\underset{j\in g_{k}}{\min}\,P(E_{1}\mid E_{2},Z_{k}=j).$ (6)
Equation (5) describes the maximum discrimination obtainable between two groups. An alternative and equivalent formulation is given in Equation (6[7]. The latter is computationally more efficient as it requires $\mathcal{O}(2|g|)$ operations compared to $\mathcal{O}(|g|^{2})$ operations for the former.
A more general approach to measure discrimination is to calculate some form of correlation coefficient between the protected attribute and the outcome. The correlation coefficient can be calculated using Pearson’s correlation [22], Spearman or Kendall’s rank correlation [23, 24]. The discrimination measure can then be defined as the absolute value of the correlation coefficient:
$\psi(\mathcal{D})=|\text{Corr}(E_{1},Z_{k})|.$ (7)
This approach can be applied to any number of groups. Fairlearn provides a pre-processing method that removes the correlation between the protected attribute and the outcome by transforming the data [7]. However, the given approach violates data integrity constraints as categorical attributes are transformed into continuous values. Further, zero correlation does not imply independency between two variables.
3.2 Multiple Protected Attributes $(|Z|>1)$
There are several ways to measure discrimination for multiple protected attributes ($|Z|>1$). Based on the works of Kearns et al. [20] and Kang et al. [12], we categorize them into two approaches: intersectional and non-intersectional (see Figure 1). Intersectional approaches consider the intersection of groups/identities. The overlapping of such identities forms subgroups [20]. Non-intersectional approaches treat each protected attribute independently.
3.2.1 Intersectional Discrimination
The central idea of intersectionality is that individuals experience overlapping forms of oppression or privilege based on the combination of multiple social categories they belong to. In the following, we will introduce necessary definitions to formulate intersectional discrimination, which is based on the work of Kearns et al. [20].
Definition 3.1 (Subgroup [20]).
Let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables representing protected attributes that can take on values from corresponding groups $g_{1},\ldots,g_{p}$. A subgroup $i$ is defined as $i=(i_{1},\ldots,i_{p})\in g_{1}\times\ldots\times g_{p}$. In other words, a subgroup encompasses multiple groups from different protected attributes.
Definition 3.2 (Subgroup Treatment).
Let $i$ be a subgroup as defined in Definition 3.1 and let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Subgroup treatment is then defined as:
$P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p}).$
Definition 3.3 (Subgroup Disparity).
Let $Z=\{Z_{1},\ldots,Z_{p}\}$ be a set of discrete random variables. Let $i,j\in g_{1}\times\ldots\times g_{p}$ be two subgroups with $i=(i_{1},\ldots,i_{p})$ and $j=(j_{1},\ldots,j_{p})$. The disparity between two subgroups is denoted as $\hat{\delta}_{Z}$ and is given by:
$\hat{\delta}_{Z}(i,j,E_{1},E_{2})=|P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p})|.$
Similarly to Equation (4), we can calculate the discrimination score for multiple protected attributes by aggregating disparities across all subgroups. A subgroup can be treated like a normal group. According to Definition 3.1, there are theoretically at least $2^{p}$ subgroups, where $p$ is the number of protected attributes. However, not all subgroups may be available in the dataset. For unavailable subgroups, the disparity cannot be calculated as the corresponding treatment is undefined.
Let denote the set of available subgroups as $G_{\text{avail}}\subseteq g_{1}\times\ldots\times g_{k}$. To finally capture the discrepancies across all available subgroup pairs, an aggregation function $\operatorname{agg}^{(1)}$ is applied to the subgroup disparities $\delta_{Z}$:
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\operatorname{agg}^{(1)}}\hat{\delta}_{Z}(i,j,E_{1},E_{2}).$ (8)
Equation (8) represents the aggregated discrimination between all available subgroups in the dataset. When using the maximum function as the aggregator, the calculations are equivalent to Equation (5) and Equation (6). The only difference is that the conditionals are now subgroups instead of groups:
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{Z_{k}}(i,j,E_{1},E_{2})$ (9)
$=\underset{i\in G_{\text{avail}}}{\max}\,P(E_{1}\mid E_{2},Z_{1}=i_{1},\ldots,Z_{p}=i_{p})-\underset{j\in G_{\text{avail}}}{\min}\,P(E_{1}\mid E_{2},Z_{1}=j_{1},\ldots,Z_{p}=j_{p}).$ (10)
Kang et al. [12] also dealt with intersectional discrimination in their work by introducing a multivariate random variable where each dimension represents a protected attribute. Their fairness objective is to minimize the mutual information between the outcome and the multivariate random variable. By minimizing the mutual information, the outcome is independent of the protected attributes, which is a desirable property for fairness [13, 25]. However, this approach relies on expensive techniques to approximate the mutual information. Using our notation, their formulation can be written as [12]:
$\psi_{\text{MI}}(\mathcal{D})=\text{MI}(E_{1},Z),$ (11)
where MI denotes the mutual information and $Z$ is the multivariate random variable representing the protected attributes. We can trivially conclude that zero mutual information implies zero intersectional discrimination [12].
Table 1: Example dataset of individuals receiving a favorable ($Y=1$) or unfavorable ($Y=0$) outcome. The dataset shows four individuals with their respective age group and sex.
Individual Age Sex Outcome ($Y$)
1 Old Male 1
2 Old Female 0
3 Young Male 0
4 Young Female 1
3.2.2 Non-intersectional Discrimination
The problem with intersectional approaches is that they can lead to an upwards bias in the discrimination score when using meta-metrics [21]. In our case, the number of subgroups grows exponentially with the number of protected attributes. This leads to many subgroups where the number of samples in each subgroup is small, resulting in larger noise in the treatment estimates [21].
Hence, we present a non-intersectional approach, which is more suitable when dealing with a large number of subgroups or when intersectional discrimination is not deemed important. The non-intersectional approach treats each protected attribute independently and aggregates the discrimination scores across all protected attributes. For this, a second aggregate function with $\operatorname{agg}^{(2)}$ is introduced, yielding the following equation:
$\psi_{\text{multi.}}(\mathcal{D})=\underset{Z_{k}\in Z}{\operatorname{agg}^{(2)}}\left\{\underset{i,j\in g_{k},i<j}{\operatorname{agg}^{(1)}}\delta_{Z_{k}}(i,j,E_{1},E_{2})\right\}.$ (12)
The first-level aggregator $\operatorname{agg}^{(1)}$ aggregates disparities within a protected attribute, considering unique pairs of groups $i$ and $j$ in each attribute. The second-level aggregator $\operatorname{agg}^{(2)}$ then combines the results across all protected attributes. By applying both operators, we obtain a discrimination measure that captures disparities between groups across multiple attributes.
3.2.3 Example
[…]
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{Z}(i,j,(Y=1),\text{True})$ (13)
$=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{\text{\{Age, Sex\}}}(i,j,(Y=1),\text{True})$ (14)
$=\underset{i\in G_{\text{avail}}}{\max}P(Y=1\mid Z_{1}=i_{1},Z_{2}=i_{2})-\underset{j\in G_{\text{avail}}}{\min}P(Y=1\mid Z_{1}=j_{1},Z_{2}=j_{2})$ (15)
$=|P(Y=1\mid\text{Age}=\text{Old},\text{Sex}=\text{Male})-P(Y=1\mid\text{Age}=\text{Young},\text{Sex}=\text{Male})|=1.$ (16)
We show in the following that the discrimination for the non-intersectional approach is different as it treats each protected attribute independently:
$\psi_{\text{multi.}}(\mathcal{D})$ $=\underset{Z_{k}\in Z}{\max}\left\{\underset{i,j\in g_{k},i<j}{\max}\delta_{Z_{k}}(i,j,(Y=1),\text{True})\right\}$ (17)
$=\max\left\{\delta_{\text{Age}}(\text{Old},\text{Young},(Y=1),\text{True}),\delta_{\text{Sex}}(\text{Male},\text{Female},(Y=1),\text{True})\right\}$ (18)
$=\max\{|0.5-0.5|,|0.5-0.5|\}=\max\{0,0\}=0.$ (19)
The non-intersectional approach yields a discrimination score of 0 because the disparities for both protected attributes are 0. This is quite different from the intersectional approach, which reports a discrimination score of 1. As seen, the choice of the approach can lead to different results.
4 Experiments
Our experimentation follows a pipeline consisting of data pre-processing, bias mitigation, model training, and evaluation. To mitigate bias in tabular datasets with multiple protected attributes, we used a sampling method, FairDo [9], that constructs fair datasets by selectively sampling data points. The method is very flexible and only requires the user to define the discrimination measure that should be minimized. In our case, we are interested in yielding a fair dataset that minimizes the bias for multiple protected attributes. The experiments revolve around two research questions:
  • RQ1 Is it possible to yield a fair dataset with FairDo, where bias for multiple protected attributes is reduced?
  • RQ2 Are machine learning models trained on fair datasets more fair in their predictions than those trained on original datasets?
4.1 Experimental Setup
Datasets and Pre-processing
Table 2: Overview of Datasets
Dataset Samples Feats. Label Protected Attributes Description
Adult [14] 32 561 21 Income Race: White, Black, Asian-Pacific-Islander, American-Indian-Eskimo, Other
Sex: Male, Female
Indicates individuals earning over \$50,000 annually
Bank [15] 41 188 50 Term deposit
subscription
Job: Admin, Blue-Collar, Technician, Services, Management, Retired, Entrepreneur, Self-Employed, Housemaid, Unemployed, Student, Unknown
Marital Status: Divorced, Married, Single, Unknown
Shows whether the client has subscribed to a term deposit.
Compas [16] 7 214 13 2-year
recidivism
Race: African-American, Caucasian, Hispanic, Other, Asian, Native American
Sex: Male, Female
Age Category: <25, 25-45, >45
Displays individuals that were rearrested for a new crime within 2 years after initial arrest.
The tabular datasets used in our experiments are: Adult [14], Bank [15], and Compas [16]. They are known for their use in fairness research and contain multiple protected attributes. We pre-processed the datasets by applying one-hot encoding to categorical variables and label encoding to protected attributes. Table 2 shows important characteristics of the datasets after pre-processing.
Each dataset was divided into training and testing sets using an 80/20 split, respectively. We ensured that the split was stratified (if possible) based on protected attributes to maintain representativeness across different groups in both sets.
Bias Mitigation
Applying the bias mitigation method FairDo [9] to the datasets can be regarded as a pre-processing step, too. This is because the method simply returns a dataset that is fair with respect to the given discrimination measure. FairDo [9] offers a variety of methods to mitigate bias and we chose the undersampling method that removes samples. The fairness objective is stated as [9]:
$\min_{\mathcal{D}_{\text{fair}}\subseteq\mathcal{D}}\quad\psi(\mathcal{D}_{\text{fair}}),$ (20)
where $\mathcal{D}$ is the training set of Adult, Bank, or Compas, and $\psi$ is the fairness objective function. We considered $\psi_{\text{intersect.}}$ and $\psi_{\text{multi.}}$ as our objectives. The first measures intersectional discrimination, while the latter measures discrimination for multiple protected attributes where the intersection of groups is not considered. Bias mitigation is only applied to the training set and the testing set remains unchanged. FairDo internally uses genetic algorithms for optimization. We used the same settings and operators as provided in the package and only adjusted the population size (200) and the number of generations (400).
Model Training
We utilized the scikit-learn library [26] to train various machine learning classifiers, namely Logistic Regression (LR), Support Vector Machines (SVM), Random Forest (RF), and Artificial Neural Networks (ANNs). These classifiers were trained on both the original and fair datasets. Classifiers trained on the original datasets serve as a baseline for comparison. We used the default hyperparameters given by scikit-learn for each classifier.
Evaluation Metrics
We evaluated the models’ predictions on fairness and performance using the test set. For fairness, we assessed $\psi_{\text{intersect.}}$ and $\psi_{\text{multi.}}$. For the classifiers’ performances, we report the area under the receiver operating characteristic curve (AUROC) [27], where higher values indicate better performance. Because removing data points can compromise the overall quality of the data, we also report the number of subgroups before and after bias mitigation, as well as the number of samples.
Trials
For each dataset, the bias mitigation process was repeated 10 times. The results were averaged over the trials to obtain a more robust evaluation.
4.2 Results
Fair Dataset Generation
Table 3: Average discrimination and number of subgroups before and after pre-processing the training sets with FairDo.
Disc. Before Disc. After Subgroups Before Subgroups After
Dataset Metrics
Adult $\psi_{\text{multi.}}$ $20\%$ $13\%$ 10 10
$\psi_{\text{intersect.}}$ $31\%$ $16\%$ 10 10
Bank $\psi_{\text{multi.}}$ $24\%$ $5\%$ 48 48
$\psi_{\text{intersect.}}$ $33\%$ $15\%$ 48 46.2
Compas $\psi_{\text{multi.}}$ $30\%$ $5\%$ 34 34
$\psi_{\text{intersect.}}$ $100\%$ $17\%$ 34 28.8
Table 3 shows the average discrimination before and after mitigating bias in the training sets. On all datasets, discrimination was reduced after applying FairDo. It can be observed that calculating the discrimination for multiple protected attributes while considering intersectionality leads to a higher discrimination value. Without considering group intersections, discrimination was reduced by 7%, 19%, and 25% for Adult, Bank, and Compas, respectively. When considering intersectionality, the discrimination was reduced by 15%, 18%, and 83%. Hence, discrimination was reduced by 28% on average across all datasets.
We also report the number of subgroups before and after bias mitigation to assess the impact of the undersampling method on the dataset. The removal of subgroups can only be observed in the intersectional setting. In the Compas dataset 5.2 out of 34 subgroups were removed on average, indicating the largest amount of subgroups removed across all datasets. While the Bank dataset consists of 48 subgroups, only 1.8 subgroups were removed on average. Because the Compas dataset’s initial intersectional discrimination score is 100%, removing multiple subgroups seems inevitable to reduce discrimination.
Model Performance and Fairness
(a) Adult (LR)
(b) Adult (SVM)
(c) Adult (RF)
(d) Adult (ANN)
(e) Bank (LR)
(f) Bank (SVM)
(g) Bank (RF)
(h) Bank (ANN)
(i) Compas (LR)
(j) Compas (SVM)
(k) Compas (RF)
(l) Compas (ANN)
Figure 2: Results on the test set. The x-axis represents the discrimination values (legend indicates used measure) and the y-axis represents the classifiers’ performances. We compare the pre-processed (fair) data with the original data. The points represent the average and the error bars show the standard deviation of the AUROC and discrimination values over 10 trials.
Figure 2 shows the results of the classifiers’ performances on the test set. The classifiers’ performances are displayed on the y-axis, while the discrimination values are shown on the x-axis. We show AUROC to evaluate the classifiers’ performances, where higher values indicate better performances. We note that the axes are not shared across the subfigures for better visibility.
Classifiers trained on fair datasets did not suffer a significant decline of performance than those trained on original datasets. In all cases, only a slight decrease of 1%-3% in performance can be observed. This indicates that the bias mitigation process does not compromise the dataset’s fidelity and therefore, the classifiers’ performances. Regarding discrimination, a great reduction can be seen. The x-axis scales are much larger than the y-axis scales, suggesting that changes in discrimination are larger than changes in performance. For example, the RF classifier trained on the Bank dataset (Figure 2g) shows a decrease in intersectional discrimination from 38% to 15%, while the performance only decreases by 2%. The results suggest that FairDo can be reliably used to mitigate bias in tabular datasets for various measures that consider multiple protected attributes. Still, we advise users to carefully perform similar analyses when applying the method to their datasets.
5 Discussion
The results of our experiments show that the presented measures detect discrimination in datasets with multiple protected attributes differently. When using the intersectional discrimination measure, more groups are identified and compared to each other, which leads to a higher reported discrimination score. While subgroups are not ignored by this measure, exponentially more groups are considered, and measuring higher discrimination scores by random chance becomes more likely [20, 21]. In contrast, treating each protected attribute separately prevents this issue but may lead to an underestimation of discrimination. The choice of measure is up to the stakeholders and depends on the context of the dataset. It also depends on the regulations for the application of the AI system. We generally recommend using the intersectional discrimination measure if the interplay of protected attributes is of interest and the number of groups is low. If a higher number of groups is present, the number of individuals in each subgroup should be high enough to prevent high discrimination scores that are due to noise. Otherwise, treating each protected attribute separately is more suitable.
Using the mitigation strategy FairDo [9], the resulting datasets in the experiments have improved statistical properties regarding fairness. Whether intersectionality was considered or not, reducing discrimination in datasets was possible, which led to training fairer machine learning models without compromising their performances substantially. At the current state, the AI Act [1] does not explicitly mention intersectional discrimination nor how to deal with multiple protected attributes generally. While recital (67) states that datasets “should […] have the appropriate statistical properties”, it does not specify what these properties are. Hence, our work serves as an initial guideline for what these properties could be and how to achieve them in practice.
6 Conclusion
Datasets often come with multiple protected attributes, which makes measuring and mitigating discrimination more challenging. Most existing studies only deal with a single protected attribute, and works that consider multiple protected attributes often focus on intersectionality. In opposition to this, we have proposed a new measure that treats each protected attribute separately. This is more suitable when the number of subgroups is too high and the number of individuals in each subgroup is low. We used both measures in our experiments and used them as objectives for the bias mitigation method FairDo. On average, the disparity between protected groups was reduced by 28% in the bias-mitigated datasets compared to the original datasets. Machine learning models trained on these datasets also showed a reduction in discrimination while maintaining their performance.